# Eternal Universes

### v1

@@ C.4 A Useful Constant
The question of notation nonwithstanding, Eqs. 3.51, 3.52, Eqs. 3.53, 3.54, and / or Eqs. 3.48, 3.49 form a system of coupled ODEs which, at least upon a cursory inspection, appear formidable. However, they can be simplified with the introduction of a geodesic dependent constant.
@@
$\xi^{2}_{t^{\ell}\oplus x^{k}}=f^{4}\left(t^{\ell}\right)\tilde{g}_{F}\left(\frac{dx^{k}}{d\lambda}\frac{\partial}{\partial x^{k}},\frac{dx^{k}}{d\lambda}\frac{\partial}{\partial x^{k}}\right)$ (3.55)
@@
$\frac{d\xi^{2}_{t^{\ell}\oplus x^{k}}}{d\lambda}$ =?${=}\frac{d}{d\lambda}\left[f^{4}\left(t^{\ell}\right)\tilde{g}_{F}\left(\frac{dx^{k}}{d\lambda}\frac{\partial}{\partial x^{k}},\frac{dx^{k}}{d\lambda}\frac{\partial}{\partial x^{k}}\right)\right]$
@@
where Eq. 3.49 was inserted to evaluate $\nabla_{\frac{d}{d\lambda}}\frac{dx^{k}}{d\lambda}\frac{\partial}{\partial x^{k}}$. ∎
Utilizing $\xi^{2}_{t^{\ell}\oplus x^{k}}$ one can revisit Eq. 3.48. The potentially complex - even before integration! - term $f\left(t^{\ell}\right)\tilde{g}_{F}\left(\frac{dx^{k}}{d\lambda}\frac{\partial}{\partial x^{k}},\frac{dx^{k}}{d\lambda}\frac{\partial}{\partial x^{k}}\right)$ can be eliminated for $\frac{\xi^{2}_{t^{\ell}\oplus x^{k}}}{f^{3}\left(t^{\ell}\right)}$. The $\nabla_{\frac{\partial}{\partial t}}\frac{\partial}{\partial t}$ equation now reads
$\nabla_{\frac{\partial}{\partial t}}\frac{\partial}{\partial t}=\frac{\xi^{2}_{t^{\ell}\oplus x^{k}}}{f^{3}\left(t^{\ell}\right)}\mathrm{grad}_{\tilde{g}_{B}}f$ (3.56)

### v2

@@ C.4 A Useful Constant
The question of notation nonwithstanding, Eqs. C.61, C.62, Eqs. C.63, C.64, and / or Eqs. C.58, C.59 form a system of coupled ODEs which, at least upon a cursory inspection, appear formidable. However, they can be simplified with the introduction of a geodesic dependent constant.
@@
$\xi^{2}_{t^{\ell}\oplus x^{k}}=f^{4}\left(t^{\ell}\right)\tilde{g}_{F}\left(\frac{dx^{k}}{d\lambda}\frac{\partial}{\partial x^{k}},\frac{dx^{k}}{d\lambda}\frac{\partial}{\partial x^{k}}\right)$ (C.65)
@@
$\frac{d\xi^{2}_{t^{\ell}\oplus x^{k}}}{d\lambda}$ $\overset{?}{=}\frac{d}{d\lambda}\left[f^{4}\left(t^{\ell}\right)\tilde{g}_{F}\left(\frac{dx^{k}}{d\lambda}\frac{\partial}{\partial x^{k}},\frac{dx^{k}}{d\lambda}\frac{\partial}{\partial x^{k}}\right)\right]$
@@
where Eq. C.59 was inserted to evaluate $\nabla_{\frac{d}{d\lambda}}\frac{dx^{k}}{d\lambda}\frac{\partial}{\partial x^{k}}$. ∎
Utilizing $\xi^{2}_{t^{\ell}\oplus x^{k}}$ one can revisit Eq. C.58. The potentially complex - even before integration! - term $f\left(t^{\ell}\right)\tilde{g}_{F}\left(\frac{dx^{k}}{d\lambda}\frac{\partial}{\partial x^{k}},\frac{dx^{k}}{d\lambda}\frac{\partial}{\partial x^{k}}\right)$ can be eliminated for $\frac{\xi^{2}_{t^{\ell}\oplus x^{k}}}{f^{3}\left(t^{\ell}\right)}$. The $\nabla_{\frac{\partial}{\partial t}}\frac{\partial}{\partial t}$ equation now reads
$\nabla_{\frac{\partial}{\partial t}}\frac{\partial}{\partial t}=\frac{\xi^{2}_{t^{\ell}\oplus x^{k}}}{f^{3}\left(t^{\ell}\right)}\mathrm{grad}_{\tilde{g}_{B}}f$ (C.66)